OpenStudy (anonymous):

lim h->0 [(3+5)^5 - 3^5]/h

OpenStudy (anonymous):

@P0sitr0n

OpenStudy (p0sitr0n):

is it (3+5)^5 or (3+h)^ 5?

OpenStudy (anonymous):

(3+h)^5

OpenStudy (p0sitr0n):

ok, its derivative of f(x) = x^5 ? x=3 f'(x) = 5x^4, 5*3^4 = 405

OpenStudy (p0sitr0n):

or if you wanted to use the limit process, well youll need the Pascal triangle to expand the 5th power

OpenStudy (anonymous):

can you help me with another one?